OpenStudy (anonymous):

what mass of oxygen gas is required to make 28g of nitrogen monoxide 4NH3+5O2 --> 6H2O+4NO

OpenStudy (anonymous):

answrr is 37g but I don't know how to get it PLEASE HELP

OpenStudy (aaronq):

first you need to convert the mass you have to moles, n, because you always work in moles. moles=mass/molar mass (1) next you use the stoichiometric coefficients to build relationship, like this: \[\frac{ n _{NO} }{ 4 }=\frac{ n _{O _{2}} }{ 5 } \] once you find the moles of O2, convert back to mass using the same formula (1)

OpenStudy (anonymous):

I didn't get it. I did 28g/30.006 = 0.933 moles of NO then times by 5 then divided by 4. Which gave me 1.16625. I then times by 15.999 and it gave me 18.7 not 37g

OpenStudy (anonymous):

what am I doing wrong

OpenStudy (aaronq):

the molar mass of O2 is 32 not 16, you're only taking one oxygen atom into account

OpenStudy (anonymous):

Got it ! Thank you soooooo much! :)

OpenStudy (aaronq):

no problem! always double check your molar masses and constants

OpenStudy (anonymous):

yes. 0.933 x 5/4 which gave me 1.16 then that times 2(15.999) gave me 37 g :)

OpenStudy (anonymous):

when doing ratio for example 2NO-->5O2 would it be 2:5 or 2:10 ?

OpenStudy (aaronq):

2 to 5, because you always use the coefficients. but make sure to use the correct molar mass, molar mass of O2 is that of 2 oxygen atoms, people always forget that.

OpenStudy (anonymous):

yeah! okay thank you so much you helped alot ! :)

OpenStudy (aaronq):

no problem at allll, glad i could !